Jess's age is six years less than three times Ethan's age. The product of their ages is 45. What is each of their ages

Answer by mananth(16946)


Jess's age is six years less than three times Ethan's age. The product of their ages is 45. What is each of their ages
Let Ethan's age =x
Jess's age 3x-6
x(3x-6)= 45
3x^2-6x-45=0
divide by 3
x^2-2x-15=0
(x-5)(x+3)=0
x=5 or -3
Ignore negative
Ethan's age =5
Jess's age 9